# UrbanVLP: Multi-Granularity Vision-Language Pretraining for Urban Region Profiling

### v2

@@ Appendix A Formulation
Definition 1 (Urban Region) We follow previous research [73, 74] and evenly divide the target area (e.g. a city) into $L$ urban regions.
@@
Definition 2 (Urban Imagery) Diverse visual images, including satellite and street-view images, facilitate an intuitive analysis of urban regions. Specifically, satellite images provide a coarse-grained overview of the structural composition of a region. Each input satellite image w.r.t the urban area $g$ can be represented as $I^{sa}_{g}\in\mathbb{R}^{H\times W\times 3}$, where $H$ and $W$ denote the length and width. Street-view images are captured by vehicles from map providers during urban traversals, offering detailed visual information on the fine-grained aspects within a region. By querying the latitude and longitude range of satellite images, multiple street-view images are associated together w.r.t the urban region $g$ based on their locations $\mathcal{L}_{g}$. Each individual street-view image is represented as $I^{{sv}}_{g}\in\mathbb{R}^{H\times W\times 3}$.
@@
Definition 3 (Text Description) Textual descriptions of urban area $g$ (including satellite text $T^{sa}_{g}$ and street-view text $T^{sv}_{g}$) provide precise means for urban area analysis. Such text can be manually generated or produced using image annotation tools. Particularly, leveraging advanced LLMs enables detailed and insightful analytical descriptions of given area images. Notably, incorporating spatial context (e.g. vegetation coverage and POI) significantly reflects the functional characteristics of the region.
@@
Definition 4 (Urban Indicators) Urban indicators assess various aspects of a region, including population, economy, society, and public services, from multiple perspectives. A set of $L$ urban regions with $K$ indicators is represented as $\mathbf{Y}\in\mathbb{R}^{L\times K}$. In this paper, we comprehensively use population (count), GDP (in million Chinese Yuan), carbon emissions (in tons), night lights (intensity), house prices (in Chinese Yuan), and points of interest, i.ePOI, (count for different categories) as ground-truth urban indicators.
Appendix B More Details about Datasets & Tasks
B.1 CityScape Dataset
@@
We construct the $\mathtt{CityScape}$ Dataset using street-view geo-tagged images collected from Baidu Map API and generated text descriptions by LMM.
@@
  • Street-View Panoramic Image Collection. The Street-View image sampling process utilized the Baidu Maps API [^1] https://lbsyun.baidu.com/, facilitating the extraction of road network data within our predefined sampling area. Using the ArcGIS [^2] https://www.esri.com/en-us/arcgis/products/develop-with-arcgis/overview tool, we exported the road network data and established a sampling interval of 500 meters, which allowed for the comprehensive collection of street-view images at every half-kilometer point along the road network, ensuring thorough coverage that complements the satellite imagery’s scope. Following the road network sampling, panoramic street-view images were collected based on the sampling points. These images, with a resolution of 2048x664, also far exceed the 256x256 resolution of satellite imagery [74], ensuring clarity and ample details for further analysis.
@@
    As is shown in Table 2, a total of 47,582 street-level images were collected, whose data volume is 2.7 times larger than that of the satellite dataset, capturing the intricacies of the urban landscape and providing a street-level view that enhances the overall urban analysis.
    In figure 11, we showcase our sampling process for the road network, which is used to collect street-view imagery at strategically selected locations. We initially acquire road network data from the online open-source website and subsequently import it into ArcGIS with a sampling interval of 500 meters. Following this, we gather the latitude and longitude coordinates of the sampling points within our dataset area. These coordinates are then used to invoke the Baidu API to collect street-view imagery. Additionally, not all sampled points have corresponding street view images, thereby resulting in a final count of street view images that is less than the number of sampled points.
@@
  • Corresponding Text Description Generation. The textual data corresponding to each street-view image is generated from ShareGPT4V [11], which is recognized for its superior text generation capabilities compared to other contemporary open-source LMM models [48, 14, 77, 20, 10, 11], offering comprehensive and high-quality textual information. The quality of the generated data is guaranteed by our proposed PerceptionScore. We filter out the texts with PerceptionScore lower than 0.5.
@@
Figure 11: Road Network Sampling for Street-View Images on Beijing.
B.2 Downstream Dataset & Tasks
In our research, the Downstream Dataset plays a crucial role in gauging the practical implications of our multimodal framework. We collect six representative urban indicators: Carbon Emission, Population, GDP, Night Light, House Price, and POI. Each indicator illuminates a significant dimension of urban vitality and sustainability. In this section, we provide the detailed introduction and data source of our downstream benchmark.
  • Carbon Emissions: Extracted from the Open-source Data Inventory for Anthropogenic CO2 (ODIAC), this dataset offers granular carbon emission figures in 2022 that are in alignment with the spatial resolution of our satellite imagery, which is one square kilometer per image. The emissions are quantified in tons.
@@ B.2 Downstream Dataset & Tasks
  • GDP: This dataset encapsulates China’s economic development by incorporating Gross Domestic Product (GDP) figures, thereby infusing an economic dimension into our analysis of urban regions. The unit is million Chinese Yuan.
  • Night Light: Nighttime light can, to some extent, characterize the intensity of human activities, which is of great significance for urban development studies. we take the nightlight data from [87] in 2020.
  • House Price: The price of housing is is subject to the influence of numerous socio-economic factors. By observing fluctuations in housing prices, it is possible to discern, to a certain extent, the state of economic development. We collect data from Anjuke (https://www.anjuke.com) in March 2023 with yuan/$m^{2}$ unit. Individual house price represents a locational attribute. To extrapolate a regional-level characteristic, we take the average of housing prices across an urban region.
  • POI: The spatial distribution of various types of POIs (Points of Interest) within urban areas is a potent social indicator, as it reflects the underlying social and economic architecture, as well as the population’s activities and urban lifestyles. Consequently, to glean insights into these urban dynamics, we collect data of POIs from web sources, encompassing both their locations and categories. The collected data has 14 categories: Dining and Cuisine, Leisure and Entertainment, Sports and Fitness, Business and Residential, Healthcare, Financial Institutions, Tourist Attractions, Lifestyle Services, Shopping and Consumption, Automobile Related, Hotel Accommodation, Transport Facilities, Science, Education and Culture, Companies and Enterprises. To capture the unique attributes of each area, we aggregate the number of POIs by category within each region.
We construct this benchmark with the aim of conducting a thorough evaluation of the urban region representations that have been derived through our learning framework. We uniformly partition the downstream dataset into train:val:test=7:1:2. These indicators constitute the downstream task dataset, covering economic (GDP, Night Light, House Price), social (Population, POI), and environmental (Carbon Emission) indicators. The data preprocessing setting adheres to [74] with refinement. For some indicators data with relatively small differences in values, we have taken the logarithm of these values to facilitate training.
B.3 Downstream Dataset Analysis
@@
Here we visualize the distribution of some downstream socio-economic indicator data on Beijing in Figure 12. The distribution of different indicators exhibits similar trends, indicating that Beijing has a clear ’center-periphery’ structure, with the core area being densely populated and economically developed, while the peripheral areas are relatively sparse.
Beyond geographical spatial differences, the concentration trends and variability of the data are critical perspectives that merit attention. In Figure 14 we demonstrate these patterns by illustrating the normalized values of different socio-economic indicators in Beijing. Outliers with extreme maximum and minimum values will be filtered out during our processing. As we can see, the values of Carbon Emission and NightLight are relatively concentrated, showing minimal variability. The distribution of GDP values is more widespread, while Population distribution exhibits a bimodal characteristic, reflecting the inherent diversity.
@@
Figure 14: PerceptionScore distribution on $\mathtt{CityScape}$ dataset for 4 representative cities.
Appendix C In-depth Overview of Baseline Methods
The single granularity imagery-based urban region representation learning methods include:
@@ Appendix C In-depth Overview of Baseline Methods
  • ViT [15]: ViT is a successful paradigm of applying transformers to computer vision, where images are segmented into patches for fixed-length modeling. When sufficient data is available for pre-training, ViT’s performance can surpass that of CNNs [22]. In our experiment, we leverage ViT-B as the baseline for comparison.
@@
  • PG-SimCLR [73]: A contrastive learning framework which introduces geographic information (i.e., POI) into urban region representation learning. In our study, we do not employ PG-SimCLR to predict POI to ensure a fair comparison.
@@
  • UrbanCLIP [74]: A model that utilizes an image-text contrastive learning-based approach for learning robust representations from urban imagery, intending to capture the complexity and diversity of urban areas.
@@
  • UrbanCLIP-SV [74]: Original UrbanCLIP is pretrained on satellite image-text pairs. Here we also verify its performance on our proposed $\mathtt{CityScape}$ dataset, utilizing the same aggregation method to acquire regional features.
We also apply multiple baselines for multi-granularity urban imagery-based representation learning:
@@
  • Vision-LSTM [27]: A vision long short-term memory neural network which could process variable-length street-view image within each region. The integration of satellite and street-view imagery, along with trajectories, facilitates a more effective capture of comprehensive urban representations, yielding reliable recognition results.
@@
  • StructuralUrban[41]: The graph-based framework for urban region profiling leveraging street segments as containers to adaptively fuse the features of multi-level urban images. We term this method StructuralUrban for simplicity.
We present a visual representation of the performance in the main paper Figure 4 to effectively demonstrate the comprehensive capabilities of our model across various downstream tasks. There is also a trend that models tend to perform better on environmental (Carbon Emission) and social (Population, POI) indicators than on economic (GDP, Night Light, House Price) indicators. This is possibly because environmental and social indicators usually have more distinct geographic distribution characteristics. For instance, neighboring regions may have similar climates, vegetation cover, and population densities, which are important factors for environmental and social indicators and can be effectively captured through spatial embedding. On the other hand, economic indicators are more complex and influenced by various factors, including policies, economic activities, and market sentiments, making them harder to model. Additionally, due to the multi-class nature of POI, models may require more complex structures to understand the underlying relationships between labels, thus resulting in lower performance.
Appendix D Implementation Details
@@
In our experiment, Adam optimizer is chosen to minimize the training loss during parameter learning. A grid search on hyperparameters is conducted, where search ranges for learning rate and batch size are set as 2e-6, 2e-5, 2e-4, 2e-3, 2e-2 and 4, 8, 16, 32, 64, respectively. The data augmentation strategy follows the setting of [61]. The weight hyperparameter $\alpha$ and $\beta$ of $\mathcal{L}_{Total}$ are set to 0.5, 0.5 respectively. We run all the models on NVIDIA A800 GPUs with PyTorch.
@@
We implement the aggregation method $\operatorname{Aggr}(\cdot)$ through MLP layers. The feature fusion method f  directly utilizes additive aggregation. To ensure a fixed length for network processing, we set the maximum street-view sequence length within one region as 25, and sequences shorter than that are achieved by padding with zeros. For the semantic segmentation model to decouple visual elements with street-view images, following  [17], the architecture for the semantic model is ResNet18dilated + PPM_deepsup. Initially, the dataset comprises 150 segmentation categories. However, in line with our requirements, we narrow down the selection to 38 categories, which we further group into 13 street-view features following [17]: Person; Bike; Heavy Vehicle; Light Vehicle; Facade; Window & Opening; Road; Sidewalk; Street Furniture; Greenery - Tree; Greenery - Grass & Shrubs; Sky; Nature.
Appendix E More Details for Modality Representation

### v3

@@ A    Formulation
Definition 1 (Urban Region) We follow previous research (Xi et al. 2022; Yan et al. 2024) to evenly divide the target area (e.g. a city) into $L$ urban regions.
@@
Definition 2 (Urban Imagery) Diverse visual images, including satellite and street-view images, facilitate an intuitive analysis of urban regions. Specifically, satellite images provide a coarse-grained overview of the structural composition of a region. Each input satellite image w.r.t the urban area $g$ can be represented as $I^{st}_{g}\in\mathbb{R}^{H\times W\times 3}$, where $H$ and $W$ denote the length and width. Street-view images are captured by vehicles from map providers during urban traversals, offering detailed visual information on the fine-grained aspects within a region. By querying the latitude and longitude range of satellite images, multiple street-view images are associated together w.r.t the urban region $g$ based on their locations $\mathcal{L}_{g}$. Each individual street-view image is represented as $I^{{sv}}_{g}\in\mathbb{R}^{H\times W\times 3}$.
@@
Definition 3 (Text Description) Text descriptions of urban area $g$ (including satellite text $T^{st}_{g}$ and street-view text $T^{sv}_{g}$) provide precise tools for urban area analysis. Such texts can be manually generated or produced using image annotation. Particularly, leveraging advanced LLMs enables detailed and insightful analytical descriptions of given area images.
@@
Definition 4 (Urban Socioeconomic Indicators) Urban socioeconomic indicators assess various aspects of a region, including population, economy, society, and public services, and so on. A set of $L$ urban regions with $K$ indicators is represented as $\mathbf{Y}\in\mathbb{R}^{L\times K}$. In this paper, we comprehensively use population, GDP, carbon emissions, night lights, house prices , and points of interest, i.ePOI, as ground-truth urban indicators.
B    Detailed Related Work
Urban Socioeconomic Indicator Prediction. In literature, many studies have focused on learning task-specific region representations from various urban data, especially urban imagery due to its consistent updates and easy accessibility (Li et al. 2022d; Liu et al. 2023d; Xi et al. 2022). For example, Urban2Vec (Wang, Li, and Rajagopal 2020) integrated street-view imagery and POI data to learn neighborhood embeddings. Some contrastive learning approaches like PG-SimCLR (Xi et al. 2022) and UrbanCLIP (Yan et al. 2024) have shown success in representing urban regions through satellite images. However, the aforementioned works exclusively considered a single type of urban imagery, overlooking the potential synergy between satellite imagery and street-view imagery, which can complement each other. Recently, some efforts (Barbierato et al. 2020; Huang et al. 2023; Li et al. 2022d; Cao et al. 2023; Law, Paige, and Russell 2019; Cao et al. 2018; Huang et al. 2023) have begun to explore the simple combination of satellite imagery and street-view imagery. Vision-LSTM (Huang et al. 2023) integrates diverse street-level images into the satellite image domain using an LSTM-based network for precise urban village identification. Li et al. (2022d) propose utilizing street segments for region modeling and predicting socioeconomic indicators by merging street-view and satellite images through a multi-stage pipeline. Nevertheless, prior studies mostly overlook distinct features in satellite and street-view images for varied granularities. To this end, we propose a unified multi-granularity cross-modal alignment framework for comprehensive urban representation in this study.
Vision-Language Pretraining (VLP). VLP aims to jointly encode vision and language in a fusion model. Early works on VLP can be broadly categorized into single-stream (Chen et al. 2020; Kim, Son, and Kim 2021; Li et al. 2020) and two-stream (Tan and Bansal 2019; Lu et al. 2019) methodologies, with the former using a unified architecture for modality integration and the latter employing separate encoders before merging modalities. Milestone work CLIP (Radford et al. 2021) and its variants (Li et al. 2021; Yao et al. 2022) highlight the efficacy of contrastive learning in cross-modal downstream tasks, such as zero-shot learning and cross-modal retrieval. Recent works (Tsimpoukelli et al. 2021; Alayrac et al. 2022) shift towards leveraging LLMs knowledge for vision-language representation learning. In urban region profiling, the potential of VLP paradigm remains untapped, with limited exploration of the benefits of textual information. Our groundbreaking work introduces multiple types of text descriptions as supplements, aiming to learn more comprehensive and interpretable urban region representations.
Large Multimodal Model (LMM). As LLMs undergo rapid evolution, a faction within the research community is increasingly focused on integrating visual knowledge into these models, such as CLIP (Radford et al. 2021) and ALIGN (Jia et al. 2021). Early studies (Li et al. 2022a, 2023) enhanced CLIP with refined data strategies for more diverse datasets, which have proven effective for fundamental visual tasks (Li et al. 2022b; Liu et al. 2023c; Zhang et al. 2022) but have shown limitations for complex tasks such as visual question answering. Recently, new models like MiniGPT-4 (Chen et al. 2023a), LLaVA (Liu et al. 2023b), and InstructBLIP (Dai et al. 2023) have improved the understanding of complex problems by expanding parameters and training data, called large multimodal models (LMMs). In this paper, we comprehensively evaluate the text generation quality of representative LMMs model and propose an automated text description generation and calibration technique to ensure effective image description capabilities with LMMs.
C    More Details about Datasets & Tasks
CityView Dataset
@@ C    More Details about Datasets & Tasks
We construct the $\mathtt{CityView}$ Dataset using street-view geo-tagged images collected from Baidu Map API following previous works (Liu et al. 2023d) and generated text descriptions by LMM. We collect and generate satellite data following the practice in  (Yan et al. 2024).
@@
  • Street-View Panoramic Image Collection. The Street-View image sampling process utilized the Baidu Map API [^1] https://lbsyun.baidu.com/, after the extraction of road network data within our predefined sampling area. Using the ArcGIS [^2] https://www.esri.com/en-us/arcgis/products/develop-with-arcgis/overview tool, we first exported the road network data. We then set a sampling interval of 500 meters, allowing for the collection of street-view images at every half-kilometer point along the road network. This approach ensured thorough coverage of the corresponding area, as demonstrated in Figure 11. Based on the sampling points, we collected panoramic street-view images at a resolution of 2048x664, which significantly exceeds the 256x256 resolution of satellite imagery (Yan et al. 2024). This higher resolution ensures greater clarity and provides ample detail for further analysis.
@@
    As is shown in Table 2, a total of 47,582 street-view images were collected, whose data volume is 2.7 times larger than that of the satellite data, capturing the intricacies of the urban landscape and providing a street-level view that enhances the overall urban analysis.
@@
  • Corresponding Text Description Generation. The textual data corresponding to each street-view image is generated from ShareGPT4V (Chen et al. 2023b), which is recognized for its superior text generation capabilities compared to other contemporary open-source LMM models (Liu et al. 2023a; Dai et al. 2023; Ye et al. 2023; Gao et al. 2023; Chen et al. 2023a, b). The quality of the generated data is guaranteed by our proposed PerceptionScore. We filter out the texts with PerceptionScore lower than 0.6.
Figure 11: Road Network Sampling for Street-View Images on Beijing.
@@
Downstream Dataset & Tasks
In our research, the Downstream Dataset plays a crucial role in gauging the practical implications of our multimodal framework. We collect six representative urban socioeconomic indicators: Carbon Emission, Population, GDP, Night Light, House Price, and POI. Each indicator illuminates a significant dimension of urban vitality and sustainability. In this section, we provide the detailed introduction and data source of our downstream benchmark.
  • Carbon Emissions: Extracted from the Open-source Data Inventory for Anthropogenic CO2 (ODIAC), this dataset offers granular carbon emission figures in 2022 that are in alignment with the spatial resolution of our satellite imagery, which is one square kilometer per image. The emissions are quantified in tons on a monthly basis.
@@
  • GDP: Our dataset encapsulates China’s economic development by incorporating Gross Domestic Product (GDP) figures, thereby infusing an economic dimension into our analysis of urban regions. The unit is million Chinese Yuan.
  • Night Light: Nighttime light can, to some extent, characterize the intensity of human activities, which is of great significance for urban development studies. We take the nightlight data from (Zhong, Yan, and Li 2022) in 2020.
  • House Price: The price of housing is subject to the influence of numerous socioeconomic factors. By observing fluctuations in housing prices, it is possible to discern, to a certain extent, the state of economic development. We collect data from Anjuke (https://www.anjuke.com) in March 2023 with yuan/$m^{2}$ unit. Since individual house prices represent locational attributes, we extrapolate a regional-level characteristic by averaging housing prices across an urban region.
  • POI: The spatial distribution of various types of POIs (Points of Interest) within urban areas is a potent social indicator, as it reflects the underlying social and economic architecture, as well as the population’s activities and urban lifestyles. Consequently, to glean insights into these urban dynamics, we collect data of POIs from web sources, encompassing both their locations and categories. The collected data has 14 categories: Dining and Cuisine; Leisure and Entertainment; Sports and Fitness; Business and Residential; Healthcare; Financial Institutions; Tourist Attractions; Lifestyle Services; Shopping and Consumption; Automobile Related; Hotel Accommodation; Transport Facilities; Science, Education and Culture; Companies and Enterprises. To capture the unique attributes of each area, we aggregate the number of POIs by category within each region.
In accordance with established conventions (Liu et al. 2023d; Yan et al. 2024; Xi et al. 2022), we uniformly and randomly partition the downstream dataset into train:val:test=7:1:2. The data preprocessing setting adheres to (Yan et al. 2024) with refinement. For some indicators data with relatively small differences in values, we have taken the logarithm of these values to facilitate training.
Downstream Dataset Analysis
To provide readers with an intuitive understanding, we visualize the distribution of three downstream socioeconomic indicators for Beijing in Figure 12. The distribution of different indicators exhibits similar trends, indicating that Beijing has a clear “center-periphery” structure, with the core area being densely populated and economically developed, while the peripheral areas have relatively sparse economic and demographic data.
@@
Beyond geographical spatial differences, the concentration trends and variability of the data are critical perspectives that merit attention. In Figure 13 we demonstrate these patterns by illustrating the normalized values of different socioeconomic indicators in Beijing. Outliers with extreme maximum and minimum values will be filtered out during our pre-processing. As we can see, the values of Carbon Emission and Night Light are relatively concentrated, showing minimal variability. The distribution of GDP values is more widespread, while Population distribution exhibits a bimodal characteristic, reflecting the inherent diversity.
@@
D    Motivation Demonstration for Multi-granularity Modeling
Here, we provide a more detailed example to illustrate our motivation for urban multi-granularity modeling. Urban environments exhibit a spatial hierarchy in reality, from a macro region level to a micro location level (e.g., architectural details, street furniture, storefront). As depicted in Figure 14, while the satellite images (a) and (b) possess remarkably similar layouts, their urban indicators considerably differ. Upon zooming into micro levels using the corresponding street-view images (c) and (d), a more nuanced and fine-grained understanding becomes apparent. We can observe that their differences arise from the varying surrounding environments (i.e., educational buildings vs. business districts). Exclusively capturing macro-level patterns through satellite imagery may introduce substantial bias, falling short in furnishing more comprehensive information across diverse urban levels.
Figure 14: Single-granularity vs. Multi-granularity modeling.
E    In-depth Overview of Baseline Methods
The single granularity imagery-based USI prediction methods include:
@@ E    In-depth Overview of Baseline Methods
  • ViT (Dosovitskiy et al. 2021): ViT is a successful paradigm of applying transformers to computer vision, where images are segmented into patches for fixed-length modeling. When sufficient data is available for pre-training, ViT’s performance can surpass that of CNNs (He et al. 2016). In our experiment, we leverage ViT-B as the baseline for comparison.
@@
  • PG-SimCLR (Xi et al. 2022): A contrastive learning framework which introduces geographic information (i.e., POI) into urban region representation learning. In our study, we do not employ PG-SimCLR to predict POI to ensure a fair comparison.
@@
  • UrbanCLIP (Yan et al. 2024): A model that utilizes an image-text contrastive learning-based approach for learning robust representations from urban imagery, intending to capture the complexity and diversity of urban areas.
@@
  • UrbanCLIP-SV (Yan et al. 2024): Original UrbanCLIP is pretrained on satellite image-text pairs. Here we also verify its performance on street-view image-text pairs of our proposed $\mathtt{CityView}$ dataset, utilizing the same aggregation method to acquire regional features.
We also apply multiple baselines for multi-granularity USI prediction:
@@
  • Vision-LSTM (Huang et al. 2023): A vision long short-term memory neural network which could process variable-length street-view image within each region. The integration of satellite and street-view imagery, along with trajectories, facilitates a more effective capture of comprehensive urban representations, yielding reliable recognition results. In our experiments, we include only satellite and street-view data, without incorporating trajectory data.
@@
  • StructuralUrban (Li et al. 2022d): The graph-based framework for urban region profiling leveraging street segments as containers to adaptively fuse the features of multi-level urban images. We term this method StructuralUrban for simplicity.
F    Implementation Details
@@
In our experiment, Adam optimizer (Kingma and Ba 2015) is chosen to minimize the training loss during the parameter learning. A grid search on hyperparameters is conducted, where search ranges for learning rate and batch size are set as 2e-6, 2e-5, 2e-4, 2e-3, 2e-2 and 4, 8, 16, 32, 64, respectively. The data augmentation strategy follows the setting of (Radford et al. 2021). The weight hyperparameter $\alpha$ and $\beta$ of $\mathcal{L}_{Total}$ are set to 0.5, 0.5 respectively. We run all the models on NVIDIA A800 GPUs with PyTorch 1.13.1 on Ubuntu 22.04. The number of epochs for the pretraining and fine-tuning stages is set to 10 and 100, respectively, with an early stopping strategy.
@@
We implement the aggregation method $\operatorname{Aggr}(\cdot)$ through MLP layers. The feature fusion method f  directly utilizes additive aggregation (the experiments regarding the choice of fusion method are provided in Appendix J). To ensure a fixed length for network processing, we set the maximum street-view sequence length within one region as 25, and sequences shorter than that are achieved by padding with zeros. For the semantic segmentation model to decouple visual elements with street-view images, following the successful precedent set by  (Fan et al. 2023b), the architecture for the semantic model is ResNet18dilated + PPM_deepsup. Initially, the dataset comprises 150 segmentation categories. However, in line with our requirements, we narrow down the selection to 38 categories, which we further group into 13 street-view features following (Fan et al. 2023b): Person; Bike; Heavy Vehicle; Light Vehicle; Facade; Window & Opening; Road; Sidewalk; Street Furniture; Greenery - Tree; Greenery - Grass & Shrubs; Sky; Nature.
G    More Details for Modality Representation